OpenStudy (anonymous):

If {x+cot40/tan50} - {1/2(cos35/sin35)}=4;find x

OpenStudy (oblivious):

Easy Way: solve for x. Plug into calculator hard way: using trig identities (look up table of trig identities online, them memorize) \[\cot(40)/\tan(50) = \cos(40)\cos(50)/\sin(40)\sin(50)\] \[=(1/2(\cos(-10)+\cos(90)))/(1/2(\cos(-10)+\cos(90)))\] = 1/2 (assuming you meant degrees here) i don't see an identity solution to the second part off the top of my head, but there may be one cot(35) = 2.11 thus we have x+1/2-1.055 = 4 or x = 4.555